OpenStudy (anonymous):

Franco bought dried beans to make soup. He bought 3/5 pounds of kidney beans, 1/2 pounds of black beans, and 9/10 pounds of navy beans. How many pounds will the beans weigh when he puts them all together?

OpenStudy (shane_b):

You just need to add up the fractions:\[\frac{3}{5}lbs+\frac{1}{2}lbs+\frac{9}{10}lbs=?lbs\]Do you know how to do that?